Circular No. 6003

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994R IN IC 2627 Ralph Martin and Andrew Williams, running the Perth Astronomy

Research Group automated supernova search on the 0.61-m Perth-Lowell reflector, report their discovery on June 3.59 UT of a supernova (mag ∼ 16) located 21′′.5 west and 1′′.2 north of the center of IC 2627 (α = 11^{h}07^{m}26^{s}, δ = −23^{o}27′.3, equinox 1950.0). The new object was not visible on May 2. C. Pollas, Observatoire de la Côte d’Azur, reports that a poor image obtained on June 3.87 shows SN 1994R to be new compared with the POSS prints and at the brightness and location described above. S. Benetti, European Southern Observatory (ESO), reports that inspection of a preliminarily reduced CCD spectrum (range 370–700 nm, resolution 2 nm) obtained on June 4.15 with ESO 3.6-m telescope (+ EFOSC1), confirms that the object is a supernova of type II. Balmer lines of hydrogen, with P-Cyg profile, are superimposed on a relatively red continuum. The double structure seen in the absorption of Hα, and the weakness of Fe II and He I 587.6-nm lines, indicate an age of ∼ 2 weeks past explosion.

SUPERNOVA 1994O IN MARKARIAN 268 H. Meusinger, Thüringer Landessternwarte, Tautenburg, reports that

SN 1994O (cf. IAUC 5988) is clearly seen on three digitized Tautenburg Schmidt plates taken May 6.9, 7.9, and 9.9 UT; it is not present on a plate taken May 2.9. Determination of magnitudes is very difficult because of the superposition with the bright central part of the galaxy.

HD 98800 S. A. Stern, Southwest Research Institute; D. A. Weintraub, Vander-

bilt University; and M. C. Festou, Observatoire Midi-Pyrénées, Toulouse, report: “Bolometric measurements of the well-known infrared-excess star HD 98800 (K5 V; distance 14 pc) were made at the Institut de Radio Astronomie Millimétrique on Apr. 17–18 using the 7-channel bolometer of the Max-Planck-Institut für Radioastronomie. These observations were made at 1.3 mm with a beam size of 11′′ and a chop throw of 30′′. The derived 1.3-mm flux density in the central beam is 36 ± 7 mJy. Combined with the 0.8-mm flux density of 102 ± 10 reported by Rucinski (IAUC 5788), this result indicates a 0.8- to 1.3-mm spectral slope ∝ ν^{2.1±0.3}.”

1994 June 4 Daniel W. E. Green